Hello, is *c* the correct answer here?
Which expression represents an example of the distributive property?
(a) 2(x + 3) = 2x + 5
(b) 2(x + 3) = 2x + 6
(c) 2(x + 3) = 2x + 23
(d) 2(x + 3) = 2x + 3
If I got it wrong can you tell me which is right and why? Thank you!

Answer by Earlsdon(6294)

The "distributive property" means...multiply each term on the inside of the parentheses by the factor on the outside of the parentheses.
Example: a(b+c) = ab+ac Applying this to your problem:
2(x+3) = 2*x+2*3 = 2x+6